OpenStudy (howard-wolowitz):

Identify a transformation of the function ......... by observing the equation of the function .....................

OpenStudy (howard-wolowitz):

\[f(x)= \sqrt[3]{x}\]

OpenStudy (howard-wolowitz):

\[g(x)=\sqrt[3]{x} + 56\]

OpenStudy (howard-wolowitz):

So that the first one and the second equation!

OpenStudy (howard-wolowitz):

A. A horizontal shift 56 units to the right. B. A vertical shift 56 units downward. C. A horizontal shift 56 units to the left. D. A vertical shift 56 units upward.

OpenStudy (howard-wolowitz):

@iGreen.

OpenStudy (howard-wolowitz):

@MARC_

OpenStudy (howard-wolowitz):

@Michele_Laino

OpenStudy (howard-wolowitz):

Hi!

OpenStudy (michele_laino):

Hi!

OpenStudy (howard-wolowitz):

Is this one clear?

OpenStudy (michele_laino):

I think if I start from g(x), I have to traslate the graph of g(x) downward of 56 units in order to get the graph of f(x)

OpenStudy (howard-wolowitz):

So basically your thinking its B because you have to shift it downwrds right?

OpenStudy (michele_laino):

since I can write: \[y - 56 = \sqrt[3]{x}\] Now I perform this traslation: \[\begin{gathered} Y = y - 56 \hfill \\ X = x \hfill \\ \end{gathered} \] and I get: \[Y = \sqrt[3]{X}\] which is the equation of f(x)

OpenStudy (howard-wolowitz):

So you are right then its B

OpenStudy (michele_laino):

yes!

OpenStudy (howard-wolowitz):

because its the only one thats downward

OpenStudy (howard-wolowitz):

cool

OpenStudy (michele_laino):

please note that the origin of the new coordinate system, namely the system XY, is located at x=0, and y= 56

OpenStudy (michele_laino):

 Please note that your answer is D.

OpenStudy (howard-wolowitz):

ok no worry

OpenStudy (michele_laino):

thanks!

OpenStudy (howard-wolowitz):

thanks for telling me